40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 19 14:06:36 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_14-06-11.png
    Reply  Fri Apr 19 20:21:52 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_20-27-04.png
       Reply  Wed Apr 24 16:06:44 2019, gautam, Update, PSL, Innolight NPRO shutoff PMCrefl.png
          Reply  Thu Apr 25 01:03:29 2019, gautam, Update, PSL, MC trans is ~1000 cts (~7%) lower than usual 
          Reply  Thu Apr 25 17:31:56 2019, gautam, Update, PSL, Innolight NPRO shutoff EdwinShutoff20190425.pngEdwinShutdown_zoomOut.png
             Reply  Mon Apr 29 16:25:22 2019, gautam, Update, PSL, PSL turned on again 
                Reply  Fri May 3 12:51:58 2019, gautam, Update, PSL, PSL turned on again 
                   Reply  Mon May 6 10:51:43 2019, gautam, Update, PSL, PSL turned off again NPROshutoff.png
                   Reply  Wed May 8 19:04:20 2019, rana, Update, PSL, PSL turned on again 
                      Reply  Thu May 9 19:50:04 2019, gautam, Update, PSL, PSL turned off again Screenshot_from_2019-05-09_19-49-29.png
                         Reply  Mon May 13 10:45:38 2019, gautam, Update, PSL, PSL turned ON again IMG_7431.JPG
                            Reply  Wed May 15 10:56:47 2019, gautam, Update, PSL, PSL turned ON again 
Message ID: 14593     Entry time: Fri May 3 12:51:58 2019     In reply to: 14583     Reply to this: 14595   14597
Author: gautam 
Type: Update 
Category: PSL 
Subject: PSL turned on again 

Per instructions from Coherent, I made the some changes to the NPRO settings. The value we were operating at is in the column labelled "Operating value", while that in the Innolight test datasheet is in the rightmost column. I changed the Xtal temp and pump current to the values Innolight tested them at (but not the diode temps as they were close and they require a screwdriver to adjust), and turned the laser on again at ~1245pm local time. The acromag channels are recording the diagnostic information.

update 2:30pm - looking at the trend, I saw that D2 TGuard channel was reporting 0V. This wasn't the case before. Suspecting a loose contact, I tightened the DSub connectors at the controller and Acromag box ends. Now it too reports ~10V, which according to the manual signals normal operation. So if one sees an abrupt change in this channel in the long trend since 1245pm, that's me re-seating the connector. According to the manual, an error state would be signalled by a negative voltage at this pin, up to -12V. Also, the Innolight manual says pin 13 of the diagnostics connector is indicating the "Interlock" state, but doesn't say what the "expected" voltage should be. The newer manual Coherent sent me has pin13 listed as "Do not use".

Setting Operating value Value Innolight tested at
Diode 1 temp [C] 20.74 21.98
Diode 2 temp [C] 21.31 23.01
Xtal temp [C] 29.39 25.00
Pump current [A] 2.05

2.10

ELOG V3.1.3-